A statistics lecturer has found that the grades on the final exam are normally distributed with a mean of 64 and a standard deviation of 11.
question: if the passing grade is 50, what proportion of the class will fail?

Answer by stanbon(75887)

A statistics lecturer has found that the grades on the final exam are normally distributed with a mean of 64 and a standard deviation of 11.
question: if the passing grade is 50, what proportion of the class will fail?
---
z(50) = (50-64)/11 = -14/11
P(x < 50) = P(z < -14/11) = 0.1016
=================
Cheers,
Stan H.
=================